A cable is to be run from a power plant on one side of a river 900 meters wide to a factory

on the other side, 300 meters downstream. The cost of running the cable under the water is $5

per meter, while the cost over land is $4 per meter. What is the most economical route over

which to run the cable?


1
Expert's answer
2021-11-25T15:31:06-0500

Let "d=" the length of cable run under the water, "300-x=" the length of cable run over the water.



Then


"Cost=C(x)=5\sqrt{(900)^2+x^2}+4(300-x), 0\leq x\leq300"


Differentiate with respect to "x"


"C'(x)=\dfrac{5(2x)}{2\sqrt{(900)^2+x^2}}-4"

Find the ctitical number(s)


"C'(x)=0=>\dfrac{5(2x)}{2\sqrt{(900)^2+x^2}}-4=0"

"5x=4\sqrt{(900)^2+x^2}"

"25x^2=16(900)^2+16x^2"

"9x^2=16(900)^2"

"3x=\pm4(900)"

"x=\pm1200"

There is no crirical number on "[0, 300]."


"C(0)=5\sqrt{(900)^2+(0)^2}+4(300-0)"

"=5(900)+4(300)=5700(\$)"

"C(300)=5\sqrt{(900)^2+(300)^2}+4(300-300)"

"=5(300)\sqrt{10}=1500\sqrt{10}\approx4743.42(\$)"

The most economical route is under the water from a power plant to a factory.